Technology resources inc., sold 6900 computer workstations, a 15% increase in sales over the previous year. How many computer workstations were sold last year?
Answer by Boreal(15235)

x=last year
x+ (.15x)=6900
1.15x=6900
divide both sides by 1.15
x=6900/1.15
6000 last year.
15% of 6000 is 900.
6000+900=6900